a gas occupies 7.84cm3 at 71.8kPa and 25 degree Celsius . find its volume at STP.


1
Expert's answer
2021-10-15T02:55:18-0400

"V_1=7.84cm^3, P_1=71.8kPa, T_1=25^oC=298K , V_2?, P_2=101.325kPa, T_2=273K"

"P_1V_1T_2=P_2V_2T_1"

"71.8\times 7.84 \times273 = 101.325\times V_2 \times298"

"V_2=5.09cm^3"